Overhead Variances, Four-Variance Analysis Oerstman, Inc., uses a standard costing system and develops its overhead rates from the current annual budget. The budget is based on an expected annual output of 126,000 units requiring 504,000 direct labor hours. (Practical capacity is 524,000 hours.) Annual budgeted overhead costs total $811,440, of which $584,640 is fixed overhead. A total of 119,000 units using 502,000 direct labor hours were produced during the year. Actual variable overhead costs for the year were $261,500, and actual fixed overhead costs were $556,150. Required: 1. Compute the fixed overhead spending and volume variances. Fixed Overhead Spending Variance $ 28,490 Favorable Fixed Overhead Volume Variance $ 32,480 Unfavorable 2. Compute the variable overhead spending and efficiency variances. Do not round intermediate calculations Variable Overhead Spending Variance $ 35,600 Unfavorable Variable Overhead Efficiency Variance $ Unfavorable
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Thanks!
Oerstman, Inc. Amount $ Note
Expected annual output 126,000.00 A
Direct labor hours required 504,000.00 B
Direct labor hours per unit                4.00 C=B/A
Annual budgeted overhead costs 811,440.00 D
Annual budgeted Fixed overhead 584,640.00 E
Annual budgeted Variable overhead 226,800.00 F=D-E
Variable overhead per hour                0.45 R=F/B
Variable Overhead Variance Finished Units Input per unit Input required Rate per unit Amount $
Standard Hours allowed for Actual Output at Standard Rate      119,000.00                  4.00       476,000.00                0.45 214,200.00 G
Actual Hours of Input, at Standard Rate       502,000.00                0.45 225,900.00 H
Actual Hours of Input, at Actual Rate 261,500.00 I
Variable overhead efficiency variance (H-G)     11,700.00 Unfavorable
Variable overhead Spending variance (I-H)     35,600.00 Unfavorable
Total Variable overhead variance (I-G)     47,300.00 Unfavorable
Fixed Overhead Variance
Annual budgeted Fixed overhead 584,640.00 J
Annual budgeted (hours) 504,000.00 K
Standard Fixed overhead per hour                1.16 L=J/K
Standard Hours allowed for Actual Output 476,000.00 M
Applied fixed overhead 552,160.00 N=M*L
Fixed Overhead Volume Variance     32,480.00 Unfavorable O=J-N
Actual Fixed cost 556,150.00 P
Standard Fixed cost 584,640.00 J
Fixed Overhead Spending Variance     28,490.00 Favorable Q=J-P